OpenStudy (anonymous):

how do i factor 2x^5-15x^3-2x=O

OpenStudy (anonymous):

x(2x^4-15x^2-2)=0 so x=0 or 2x^4-15x^2-2=0 so when 2x^4-15x^2-2=0 then u can use the common binomial factorization u should get 3 answers one of them is x=0

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (theviper):

u can close this ques as u have got the answer 29 minutes ago!